# Understanding Large Language Models in Your Pockets: Performance Study on COTS Mobile Devices

### v3

@@ V-A2 Speedup with Special Machine Instructions
To further explore the potential of Arm CPUs, we test the Llama-2-7B model on top-tier SoCs and recompiled llama.cpp directly on the device with the i8mm flag to enable INT8 matrix multiplication instructions. This ensures that the machine code consists of smmla[42] and sdot[43]. The sdot instruction accelerates vector computations by processing multiple INT8 dot products simultaneously, making it well-suited for matrix-vector multiplication during the decoding. In contrast, the smmla instruction handles block matrix multiplication and delivers twice the multiply-accumulate efficiency of sdot, making it ideal for accelerating matrix operations in the prefill stage.
We note that Arm developers have proposed the prearrangement of weights in blocks to avoid pseudo-scalar operations and fully exploit the parallel capabilities of specialized instructions [44]. In the original weight layout, only a single column of weights is processed at a time, and it requires multiple dot product operations. The optimized layout, however, distributes the weight columns across multiple computational lanes, enabling parallel processing of several weight columns in a single instruction. This approach not only increases throughput but also reduces memory access by sharing input data within lanes, further enhancing inference efficiency.
@@ V-B GPU Utilization for LLM inference
We note an intriguing exception regarding performance on Mali GPUs. The Mali-G720720 Immortalis MP12, found in the Dimensity 9300 (Vivo Pad3 Pro), has better hardware parameters but has poorer performance than Adreno $750$ in Snapdragon SoC (Xiaomi14 Pro). We measure the maximum throughput for FP16 operations on devices using clpeak[45], which is a benchmarking tool intended for calculating GPU performance by executing different OpenCL kernels. The actual maximum memory bandwidth is also obtained using clpeak and presented with the theoretical maximum bandwidth in Table IV. From Table IV, we can see that the Mali-G720 MP12 exhibits $1.5$ times the throughput of the Adreno $750$ in float16 operations. However, Figure 3 shows that its prefill speed is $4.6$ times slower.
@@
The possible reason for this discrepancy comes from the GPU utilization. Figure 8 presents the average utilization and memory bandwidth during inference on Mali-G720720, Adreno $750$, and iPad Pro. It should be clarified that, on Apple devices, Load-Store utilization is reported separately as buffer reads and buffer writes, whereas on non-Apple devices we refer to the aggregate Load-Store utilization.The results reveal that MLC LLM fails to effectively utilize ALU units on both Mali and Adreno GPUs, with inefficiency being particularly severe on Mali.
During the compute-bound prefill stage, the Mali-G720’s arithmetic unit utilization averages less than $3\%$. The Adreno 750 performs better, but only reaches around $20\%$. By contrast, the Apple M1 achieves up to $92\%$. In the memory-bound decoding stage, the Mali-G720720 shows a severe memory bottleneck for the nearly saturated Load-Store units utilization. Taken together, these findings highlight that generic operator implementations are mismatched with the Mali GPU, leaving compute resources underused in the prefill stage and memory resources overwhelmed in the decode stage.
V-B1 Understand the underuse of non-Apple GPUs
@@
During the prefill stage on the Adreno $750$, cache profiling reveals L22 miss rates exceeding $40\%$, indicating frequent cache failures and heavy reliance on system memory. Prolonged L22 stalls from memory transfers substantially reduce ALU activity. On the Mali-G720720, a similar problem arises: over $80\%$ of Load-Store units reads are partial-width, resulting in numerous short memory accesses that underutilize cache capacity [46]. Moreover, the rate of all registers warp reaches nearly $100\%$, which halves the peak thread occupancy of the shader core and makes shader performance particularly sensitive to cache misses and memory stalls[47]. Taken together, the findings demonstrate that, although the prefill stage is inherently compute-intensive, suboptimal memory access patterns constitute the primary performance bottleneck in existing implementations.
@@
Figure 10: Throughput under continuous inference (CPU)
Figure 11: Prime Core Frequency
@@ V-B2 GPU Optimization Practices
General-purpose LLM implementations on mobile GPUs do not fully harness the parallelism capabilities of these devices. The results highlight a significant performance gap between different GPU architectures (e.g., Mali vs. Adreno)—for instance, Adreno GPU may achieve $5\times$ throughput than Mali GPU (which has higher computation capacity) for the same workload. To unlock the full potential of mobile GPUs, customized implementations are essential.
@@
The primary performance bottleneck on GPUs arises from suboptimal data layouts and memory access patterns. Even in the compute-bound prefill stage, frequent external memory accesses cause significant ALU idle time. Optimizations such as restructuring data layouts or applying memory coalescing can improve cache utilization, while rebalancing thread workloads and minimizing register usage can further enhance computational efficiency.
Figure 12: Throughput under continuous inference
Figure 13: Performance Sensitivity to Starting Temperature
V-C Impact of DVFS and Thermal Throttling
V-C1 Impact of DVFS on CPU
@@
Figure 11 illustrates a general decline in performance as the rounds of inference increase. Although the inference throughput for both prompt lengths is similar, the iPad exhibits a more pronounced performance drop in long-prompt tests compared to short-prompt tests. Moreover, in short-prompt tests, unlike the stable prefill speed, the decode speed on the iPad decreases during the early stages and eventually falls from the best to a level comparable with the second-highest among the other devices. Additionally, on Xiaomi 14 Pro, throughput for both prefill and decode initially rises, then decreases, and finally stabilizes at a relatively steady level, accompanied by minor oscillations around the equilibrium. In contrast, Dimensity $9300$ (Vivo Pad3 Pro) and Kirin 90009000E (Huawei Matepad12.6 Pro) exhibit only minor latency fluctuations, with a maximum decrease of approximately $10\%$. This trend aligns with the data presented in Figure 3, which shows that Xiaomi $14$ Pro experiences a more significant performance drop as prompt length increases.
@@
V-C2 Impact of CPU Thermal Throttling
As discussed in Section V-C1, only a few consecutive inference runs are sufficient to trigger CPU frequency reduction due to thermal throttling. To further examine the impact of heat accumulation over time, we extend the experiments on Xiaomi $14$ and analyze ($1$) temperature and performance variation during sustained Inference and ($2$) the impact of initial device temperature on performance.
Impact of Heat Accumulation on Sustained Inference: Figure 13 illustrates performance variations of different workloads over $20$ consecutive inference runs and the corresponding thermal buildup. Both curves exhibit an initial rapid temperature rise followed by stabilization. During the first $500$ seconds, the trends are nearly identical, starting from about $34^{\circ}\text{C}$ and climbing above $40^{\circ}\text{C}$. In the stable stage, the rate of temperature rise slows, whereas the long-prompt tests are still faster than the short-prompt tests, leading to a higher equilibrium near $42^{\circ}\text{C}$. After about $1000$ seconds, the temperature of long-prompt tests fully stabilizes, with minor oscillations around equilibrium likely reflecting stricter thermal regulation.
@@ V-C2 Impact of CPU Thermal Throttling
V-C3 Impact of DVFS on GPU
Figure 14: Throughput under continuous inference (GPU)
@@ V-C3 Impact of DVFS on GPU
V-D Model Preparation Latency
In Section V-C1, we have studied the inference performance with warm-up. However, in real-world applications, users may trigger the model sporadically, leading to scenarios where the model must perform a cold start without a warm-up, known as the cold start. Thus, the time required to generate the first token becomes critical to user experience. The process of generating the first token involves several stages: loading model weights from external storage into memory, preparing memory for variables, and executing prefill and decoding. In this section, we focus on the preparation procedure before prefill and evaluate the latency across various devices. We perform tests on seven devices and measure the time spent on preparation with and without warm-up. If the device is tested without warm-up, we reboot it to refresh RAM and cache. Results are present in Figure 15.
@@ VI Opportunities for Improving Mobile LLM Performance
Hybrid Inference to Enhance User Experience Although CPU performance appears sufficient, as shown in Section V-C2, running inference solely on CPUs leads to rapid temperature increases. Given the strong performance of the NPU and GPU during the prefill stage, a hybrid inference approach—utilizing the NPU or GPU for prefill and the CPU for decoding—is expected to enhance efficiency while delaying temperature rise and improving the overall user experience.
VII Discussion
Large Lanuage Models. There are many lightweight LLMs, with parameters ranging from 1B to 13B, that can fit in memory on most high-end mobile devices. Additionally, strategies have been developed to run larger models by offloading part of the model weights to external storage. In this paper, rather than focusing on performance differences across models, we concentrate on the hardware capabilities to identify bottlenecks that limit LLM performance and thoroughly understand how much untapped potential remains.
Most existing generative AI models are transformer-based, where the primary operations are matrix-matrix or matrix-vector multiplications. While different models may vary in terms of matrix size or the number of transformer blocks, the core operations remain highly resource-demanding matrix computations. Understanding the hardware’s capacity to handle these operations is key to optimizing LLM performance on mobile devices.
Quantization and Model Accuracy. An important aspect of evaluating a model’s capability is the accuracy of the responses generated by the LLM. However, for on-device LLM inference, the factors influencing model accuracy are primarily subject to the model architecture and the quantization of its weights. Based on this fact, MobileAIBench[10] provides a comprehensive benchmark for analyzing accuracy across various NLP tasks under different quantization schemes for both LLMs and large multimodal models (LMMs). Since our primary focus is on hardware-related performance metrics, we have chosen the most widely used quantization scheme, which offers the best balance between speed and accuracy. While lower-bit quantization algorithms may enhance inference speed, we excluded them from this study due to their significant impact on accuracy.
Multimodality Models. Driven by the growing demand for advanced AI assistants, the deployment of large multimodal models (LMMs) on mobile devices is an inevitable trend. However, we excluded LMMs from our study for two main reasons. First, current LMMs may not yet be ready for mobile deployment. For example, MobileAIBench reports that testing Llava-Phi-2[55] on the iPhone 14 resulted in a first token generation of $66.47$ seconds. It remains unclear whether this poor performance is due to hardware constraints or inefficient operator implementation. Second, since LLMs and LMMs share significant similarities in both model architecture and operator implementations, the insights gained from on-device LLM tests can likely be generalized to large multimodal models.
VIII Concluding Remarks

### v4

@@ V-A2 Speedup with Special Machine Instructions
To further explore the potential of Arm CPUs, we select three models of varying sizes and architectures: Gemma33-11B, Qwen33-44B, and Llama22-77B. We recompiled llama.cpp directly on top-tier devices with the armv8.78.7 flag to enable INT88 matrix multiplication instructions. This ensures that the machine code includes smmla[8] and other efficient Neon instructions. Figure 9 illustrates that enabling these special machine instructions significantly improves throughput, with the acceleration primarily observed in the prefill stage. While the decode stage is memory-bound, it still shows slight improvement. The Dimensity $9300$ maintains optimal performance with all big cores, but when both using $6$ cores, the Snapdragon 88 Gen33 outperforms the Dimensity $9300$. This performance advantage is likely due to the higher CPU frequency of the Snapdragon 88 Gen33. Furthermore, when both SoCs are configured with optimal core setups, the Snapdragon 88 Gen33 experiences a slightly higher performance boost from instruction acceleration compared to the Dimensity $9300$.
For the same quantization method, larger model sizes experience more significant acceleration. The Snapdragon 88 Gen33 and the Dimensity $9300$ both show the highest speedup with Llama22-77B. Specifically, in the prefill stage, speedup is more than $6\times$ with Q44_0 and about $2\times$ with Q44_K_M. For Gemma33-11B, the speedup is around $5\times$ with Q44_0 and $1.5\times$ with Q44_K_M.
Compared to Q44_K_M, Q44_0 optimizes weight layout by rearranging weights in blocks [25]. This approach eliminates pseudo-scalar operations and fully leverages the parallel capabilities of specialized instructions. While the original weight layout processes a single column at a time, the optimized layout distributes weight columns across multiple lanes, enabling parallel processing and reducing memory access, thereby enhancing inference efficiency. For the Snapdragon 8 Gen3, the presence of small cores slows down throughput, especially with Q4_0. The performance gap between 6 and 8 cores widens with Q4_0, likely due to the mismatch between the specialized acceleration instructions (paired with prearranged weights) and the architectural constraints of the efficiency cores (i.e., small cores). Interestingly, when using Q4_0 with instruction acceleration, the Snapdragon 8 Gen3’s performance with 6 big cores outperforms the Dimensity $9300$’s 8 big cores, indicating that the specific operator implementation plays a crucial role in performance.
@@ V-B GPU Utilization for LLM inference
We note an intriguing exception regarding performance on Mali GPUs. The Mali-G720720 in the Dimensity 9300 (Vivo Pad3 Pro), has better hardware specifications but poorer inference performance than Adreno $750$ (Xiaomi14 Pro) in Snapdragon SoC. We measure the maximum throughput for FP16 operations on devices using clpeak[32], which is a benchmarking tool intended for calculating GPU performance by executing different OpenCL kernels. The actual maximum memory bandwidth is also obtained using clpeak and presented with the theoretical maximum bandwidth in Table IV. From Table IV, we can see that the Mali-G720 MP12 exhibits $1.5$ times the throughput of the Adreno $750$ in float16 operations. However, the results in Figure 3 reveals a stark contradiction, as the prefill speeds of the four tested models are approximately $10$ times slower on the Mali GPU than on the Adreno $750$.
Using Llama2-7B as a representative case, we further analyzed a set of GPU performance metrics during inference to identify the root causes of this performance gap. Figure 10 presents the average utilization and memory bandwidth during inference on Mali-G720720, Adreno $750$, and Apple M11. It should be clarified that, on Apple devices, Load-Store utilization is reported separately as buffer reads and buffer writes, whereas on non-Apple devices we refer to the aggregate Load-Store utilization.
The possible reason for this discrepancy comes from the GPU utilization. Referring to Figure 3 and Figure 10, the ALU utilization during the compute-bound prefill stage demonstrates a strong positive correlation with throughput. The Mali-G720720’s arithmetic unit utilization averages less than 33%, implying that the matrix multiplication operators are almost not parallelized. The Adreno $750$ performs better, but only reaches around $20\%$. In contrast, the Apple M11 achieves up to 9292%, providing approximately an $8\times$ speedup over the Adreno $750$. In the memory-bound decoding stage, the Mali-G720720 shows a severe memory bottleneck for the nearly saturated Load-Store units utilization. Taken together, these findings highlight that generic operator implementations are mismatched with the Mali GPU, leaving compute resources underused in the prefill stage and memory resources overwhelmed in the decode stage.
Figure 10: GPU Utilization (MLC LLM with Llama22-77B)
V-B1 Understand the underuse of non-Apple GPUs
@@
During the prefill stage on the Adreno $750$, profiling reveals that ALU underutilization is mainly caused by memory access stalls. In a single layer of Llama22-77B, the final matrix multiplication kernel experiences 4604% of its cycles stalled due to memory access, with L22 cache miss rates exceeding 6060%. This indicates that the bottleneck probably results from non-contiguous memory access, which causes prolonged L22 stalls and reduces ALU activity. For other kernels, approximately 4040% of the stalls are attributed to texture memory access, indicating that the bottleneck arises from the volume of texture memory requests exceeding the available hardware bandwidth. This suggests that the inefficiencies may be caused by redundant data requests or suboptimal data sharing between threads, which prevent the kernel from effectively reusing cached data.
On the Mali-G720720, a similar problem arises: over $80\%$ of Load-Store units reads are partial-width, resulting in numerous short memory accesses that underutilize cache capacity [11]. Moreover, the rate of all registers warp reaches nearly $100\%$, which halves the peak thread occupancy of the shader core and makes shader performance particularly sensitive to cache misses and memory stalls[12]. Taken together, the findings demonstrate that, although the prefill stage is inherently compute-intensive, suboptimal memory access patterns constitute the primary performance bottleneck in existing implementations.
@@ V-B2 GPU Optimization Practices
General-purpose LLM implementations on mobile GPUs do not fully harness the parallelism capabilities of these devices. The results highlight a significant performance gap between different GPU architectures (e.g., Mali vs. Adreno)—for instance, Adreno GPU may achieve $10\times$ throughput than Mali GPU (which has higher computation capacity) for the same workload. To unlock the full potential of mobile GPUs, customized implementations are essential.
@@
The primary performance bottleneck on GPUs arises from suboptimal data layouts and memory access patterns. Even in the compute-bound prefill stage, frequent external memory accesses and overloaded texture memory requests cause significant ALU idle time. Optimizations such as restructuring data layouts or applying memory coalescing can improve cache utilization, while rebalancing thread workloads and minimizing register usage can further enhance computational efficiency.
TABLE V: Power Consumption ([Prompt, Gen] = [$64$, $128$])
Device (CPU/GPU) #Threads Prefill Decode
Tput1 Power2 Tput1 Power2
Vivo Pad3 Pro 44 $6.11$ $0.050$ $5.01$ $0.053$
(CPU w/o IA3) 66 $8.08$ $0.039$ $6.40$ $0.047$
88 $9.80$ $0.031$ $7.09$ $0.042$
Xiaomi Pad6 Pro 44 $5.55$ $0.106$ $4.11$ $0.137$
(CPU w/o IA) 66 $5.89$ $0.106$ $4.12$ $0.143$
88 $6.08$ $0.106$ $4.11$ $0.143$
Xiaomi Pad (CPU w/ IA) 66 $12.87$ $0.038$ $4.51$ $0.117$
Xiaomi Pad (GPU) 66 $46.03$ $0.013$ $7.33$ $0.058$
Huawei Matepad12.6 Pro 44 $4.28$ $0.100$ $3.56$ $0.118$
(CPU w/o IA) 66 $4.79$ $0.137$ $3.43$ $0.111$
88 $4.99$ $0.144$ $3.60$ $0.120$
11: Throughput (tokens/s); $2$: Power (mAh/token); $3$: Instruction Acceleration
V-C Battery Consumption
While inference speed remains the primary metric for ensuring high-quality user experience, power consumption is equally critical, as it directly determines the feasibility of sustaining on-device LLM inference. To comprehensively evaluate the trade-off between throughput and energy efficiency, we evaluated three representative mobile devices across distinct performance tiers with llama.cpp and Llama22-77B. The experimental results represent the mean of two independent trials, each consisting of 20 consecutive inference iterations with a 64-token prefill and 128-token generation. For precise energy measurement, we directly sample current and voltage from system files on the rooted Xiaomi device. For non-rooted devices, energy metrics were obtained from application-specific power consumption data provided by the operating system’s monitoring tools. The results are shown in Table V.
Core Configuration and Efficiency: Adjusting the number of threads influences energy-to-throughput efficiencies across different SoC architectures. For top-tier devices, such as the Vivo Pad33 Pro which utilizes an all-big-core architecture, increasing the thread count improves efficiency. As more cores are engaged, inference speed increases, and energy efficiency improves. For example, when using all cores, prefill energy consumption drops to its lowest point at 0.031 mAh per token. In contrast, high-tier devices featuring a combination of big and little cores, such as the Xiaomi Pad66 Pro, experience diminishing returns—adding more cores provides only a modest speed improvement for prefill, but power consumption remains unchanged. For devices with a weaker big-little core configuration, such as the Huawei tablet, performance gains from multi-threading come at the expense of higher power consumption, particularly during the prefill stage, exhibiting significant energy overhead.
Prefill vs. Decode Energy Profile: Energy consumption per token shows distinct patterns depending on the device’s hardware capability. Typically, the decode stage is more energy-intensive due to its sequential, auto-regressive nature, which doesn’t benefit from batch parallelization. However, on less powerful devices, this pattern reverses. The Huawei tablet is the only device tested where prefill energy exceeds decode energy. As the thread count increases from 4 to 8, prefill energy rises from 0.1mAh/token to 0.14mAh/token, while decode energy remains nearly unchanged at approximately 0.12mAh/token.
Hardware Acceleration and Energy Efficiency: Compared to standard CPU execution, specialized hardware offers transformative efficiency. On the Xiaomi Pad 6 Pro (using 6 threads), Instruction Acceleration (IA) (see Section V-A2) speeds up prefill by $2.19\times$ while reducing energy consumption by 63.8263.82%. The GPU shows even more dramatic improvements, accelerating prefill by $7.81\times$ and cutting energy consumption by $87.79\%$. Even in the less computational-intensive decoding phase, the GPU achieves a $1.78\times$ speedup while reducing energy consumption to 0.058mAh/token, a $59.75\%$ decrease. These results underscore the significant energy efficiency advantage of GPU-based inference compared to CPU execution.
Figure 12: Throughput under continuous inference (CPU)
Figure 13: Prime Core Frequency
V-D Impact of DVFS and Thermal Throttling
V-D1 Impact of DVFS on CPU
@@
Figure 13 illustrates a general decline in performance as the rounds of inference increase. Although the inference throughput for both prompt lengths is similar, the iPad exhibits a more pronounced performance drop in long-prompt tests compared to short-prompt tests. Moreover, in short-prompt tests, unlike the stable prefill speed, the decode speed on the iPad decreases during the early stages and eventually falls from the best to a level comparable with the second-highest among the non-Apple devices. Additionally, on Xiaomi 14 Pro, throughput for both prefill and decode initially rises, then decreases, and finally stabilizes at a relatively steady level, accompanied by minor oscillations around the equilibrium. In contrast, Dimensity $9300$ (Vivo Pad3 Pro) and Kirin 90009000E (Huawei Matepad12.6 Pro) exhibit only minor latency fluctuations, with a maximum decrease of approximately $10\%$. This trend aligns with the data presented in Figure 2, which shows that Xiaomi $14$ Pro experiences a more significant performance drop as prompt length increases.
@@
V-D2 Impact of CPU Thermal Throttling
As discussed in Section V-D1, only a few consecutive inference runs are sufficient to trigger CPU frequency reduction due to thermal throttling. To further examine the impact of heat accumulation over time, we extend the experiments on Xiaomi $14$ and analyze ($1$) temperature and performance variation during sustained inference and ($2$) the impact of initial device temperature on performance.
Figure 14: Throughput under continuous inference
Figure 15: Performance Sensitivity to Starting Temperature
Figure 16: Throughput under continuous inference (GPU)
Impact of Heat Accumulation on Sustained Inference: Figure 15 illustrates performance variations of different workloads over $20$ consecutive inference runs and the corresponding thermal buildup. Both curves exhibit an initial rapid temperature rise followed by stabilization. During the first $500$ seconds, the trends are nearly identical, starting from about $34^{\circ}\text{C}$ and climbing above $40^{\circ}\text{C}$. In the stable stage, the rate of temperature rise slows, but the long-prompt tests still experience a faster rate of temperature rise compared to the short-prompt tests, leading to a higher equilibrium near $42^{\circ}\text{C}$. After about $1000$ seconds, the temperature of long-prompt tests fully stabilizes, with minor oscillations around equilibrium likely reflecting stricter thermal regulation.
@@ V-D2 Impact of CPU Thermal Throttling
V-D3 Impact of DVFS on GPU
@@
V-E Model Preparation Latency
In Section V-D1, we have studied the inference performance with warm-up. However, in real-world applications, users may trigger the model sporadically, leading to scenarios where the model must perform a cold start without a warm-up, known as the cold start. Thus, the time required to generate the first token becomes critical to user experience. The process of generating the first token involves several stages: loading model weights from external storage into memory, preparing memory for variables, and executing prefill and decoding. In this section, we focus on the preparation procedure before prefill and evaluate the latency across various devices. We perform tests on seven devices and measure the time spent on preparation with and without warm-up. If the device is tested without warm-up, we reboot it to refresh RAM and cache. Results are present in Figure 17.
@@ VI Opportunities for Improving Mobile LLM Performance
Hybrid Inference to Enhance User Experience Although CPU performance appears sufficient, as shown in Section V-D2, running inference solely on CPUs leads to rapid temperature increases. Given the strong performance of the NPU and GPU during the prefill stage, a hybrid inference approach—utilizing the NPU or GPU for prefill and the CPU for decoding—is expected to enhance efficiency while delaying temperature rise and improving the overall user experience.
VII Concluding Remarks